# Human-Centered LLM-Agent User Interface: A Position Paper

### v1

@@ 1 Introduction
Large Language Models (LLMs) can be used to connect an underlying system with a user via the natural language medium, forming an LLM-powered application, as shown in Figure 1. The LLM is embedded in an LLM-in-the-loop state machine to acquire multi-modal input/output capabilities, emulate logical reasoning and planning, and use tools or operate a system Shen et al. (2024b); Liu et al. (2023); Liang et al. (2024); Tao et al. (2023). Consequently, the user is able to indirectly but effectively use the underlying system via chatting with an LLM assistant.
However, the current applications hardly address the user-interaction potential stemming from the multi-round chatting setup. Although the user can refer to the chat history, the LLM assistant rarely challenges the user or even asks to clarify the user’s intention. Instead, the LLM closely follows the user’s command, missing the golden opportunity to improve the user’s usage scheme and his/her understanding of the system/tools. That inefficacy becomes jarringly apparent when one tries to design from scratch a new complex system with an LLM interface.
We posit that the operation scope of an LLM-Agent User Interface (LAUI) is much wider than that. The interface should be more than an assistant or a butler, but instead a secretary, actively working with the user to discover emergent interaction schemes on the fly. LAUI should be proficient with the underlying system, study the user, study the user’s needs (instead of commands), reason on its own, and propose tailored interaction schemes to the user, including what modes of feedback are provided and what input is expected from the user. In the conventional way of interaction, including GUI and current LLM-powered applications Microsoft (2024), the designer has to imagine possible usage workflows given the system capabilities at design time, and the user is expected to learn the system (via tutorials, explorations, and practicing) in order to come up with a workflow for each task. In contrast, with LAUI, the user only needs to describe his/her needs and doesn’t need to deeply understand the application, and a workflow will naturally emerge as the LLM agent works with the user. We call for more research exploring the potential of LAUI.
Figure 1: The LLM agent serves as the interface between the underlying system and the user. The LLM agent together with the system forms the application.
As a concrete illustration of LAUI, we present Flute X GPT, an LLM-in-the-loop music-tutoring application consisting of an LLM agent, a prompt manager, a software system, and hardware. The application provides real-time haptic guidance via servo motors, real-time visual music-symbol feedback, real-time audio feedback, and natural language chat, all controlled by the LLM agent. This is the first time an LLM-powered interface is applied to a working system of such complexity and real-time user interactivity.
@@ 2 Flute X GPT
We describe Flute X GPT [^1] The source code is open to public at https://github.com/DoubleBlind23408929/Flute-X-GPT, a music-tutoring application using LAUI. The human user works with Flute X GPT in workshop episodes, practicing to play the flute and learning music. The application gives real-time multi-modal feedback, including haptic feedback that applies force to the user’s fingers, visual feedback displaying performance errors, audio feedback rendering the music, and natural-language speech given by a robot music teacher. The underlying software-hardware system has numerous different configurations, each leading to a different interaction workflow. Each setting (e.g., toggle certain feedback, conditions for triggering feedback) can be controlled independently, so the number of configurations grow exponentially with the number of settings. For the user, it is unrealistic to first master the complex system before using it. Even for the designer, the space of possible interaction schemes is impossible to enumerate during design time. The LLM agent steps in to bridge that gap. Via prompting, we instruct the LLM agent to be proficient with all the raw capabilities of the system. During use time, the LLM agent converses with the user to clarify what interaction workflow will benefit the user’s music learning goal the most. The LLM agent studies the user’s preferences and diagnoses the musical challenges the user is facing. When configuring the underlying system, the LLM agent uses common sense to reason about the implied consequences of the system configurations. Certain mixtures of settings have never been previously considered by any human designer, but can still emerge during use time.
@@ 2.1 User Experience
Table 1: Functions that the LLM agent can call to control Music X Machine. The description addresses the LLM agent.
The intended user experience. The user has no prior knowledge about the flute tutoring system. The only assumption about the user is that the user wants to learn the flute. It is the LLM agent’s job to adapt to the user’s current flute playing capability, other musical skills, demographics, vocabulary, patience, style of learning, etc. During the music learning workshop, the LLM agent wears the face of a robot music teacher to chat with the user. For example, the robot teacher asks the user to put on a pair of haptic gloves, and explains that they provide force feedback at each finger. The workshop then alternates between practice sessions where the user plays a segment of a song under real-time guidance and verbal dialogues between the user and the robot teacher. The user gradually experiences more and more interaction schemes that trains musicality in different ways, and expresses their ideas about using the application and music learning in general. The LLM agent uses that opportunity to study the user and steers the workshop accordingly, aiming to maximize music education effect. The user learns to treat the robot teacher as a considerate and professional music tutoring agent capable of thinking multiple steps ahead, formulating plans with the user, and explaining the plans as well as music knowledge and education principles for the user.
We present three video demos of real-human user tests, including one scripted trial and two improves trials. [^2] Demo playlist: https://youtube.com/playlist?list=PLOplnUxtqG3SKrR01tYTdrqdaClayFTxp
Video demo: Scripted trial. In this demo, all actors follow a script generated by an offline but faithfully emulated interaction between a user and the LLM agent. To re-emphasize, the LLM agent’s lines in the script are not written by humans, but is outputted by the LLM agent. Scripting removes the latency of LLM autoregressive generation. To ensure the script is concise and demonstrates a wide range of behaviors, we intervene with the script generation process. Every turn, we select one response out of 4 to 16 candidates sampled by the LLM. Seldomly, we add a temporary user-role prompt to give a short hint to the LLM agent, or edit the generated response. All the above interventions are kept to the minimum to ensure the vast majority of the agent speech is the authentic output of the LLM. See Figure 2(a) for an excerpt from the video.
@@ 2.2 System Capabilities
Haptic guidance. The hardware includes a pair of gloves with servo motors and movable finger rings. Through these gloves, the system moves the user’s fingers to help with performance motions. The haptic guidance can be configured via various settings (e.g., is the guidance sustained throughout each note or applied at each note onset? Apply guidance for each note? For incorrectly played notes? For unplayed notes?). Certain combinations of settings form configuration presets whose interaction scheme is deemed meaningful by the designer. For example, in the Force Mode preset, every note triggers a full-force guidance for each finger. In the Adaptive Mode preset, the user plays the song on his/her own, and the gloves correct the mistakes. The haptic configuration should be tuned to adapt to the user’s skill level and the song’s difficulty Zhang et al. (2019).
Visual feedback. A monitor displays the score of the current song. The user-played notes are displayed on top of the notes in the score, yielding the real-time visual Knowledge-of-Result (KR) feedback. It is a visual cue for the user to know where he/she is in terms of the pitch and helps the user internalize the score notations Chin et al. (2020). It can be toggled on or off.
@@
Song database. The practicing music materials that the system uses is processed from the POP909 dataset Wang et al. (2020); Chin and Xia (2022). It contains pieces and sections of monophonic melody lines from pop songs.
@@ 2.4 LLM in the Loop
Figure 3 illustrates the inner workings of the LLM-in-the-loop application, Flute X GPT. The LLM we use is GPT-4 Achiam et al. (2023). The Music X Machine has been described in Subsection 2.1 and 2.2. The System Principles are a passage of prompt given to the LLM at the top of the conversation that defines the role and interaction principles for the LLM agent (see Appendix A). Here are two representative excerpts.
@@
The text-to-speech (T2S) module includes Text to Speech PRO VidLab (2023) on Rapid API and FastSpeech 2 Chien et al. (2021). The speech-to-text (S2T) module is Whisper Radford et al. (2023), prompted to ignore flute sounds. The Robot is TeoTronico Suzzi (2023) who can play the piano from MIDI, lip sync according to the speech audio amplitude in real time, and make random facial expressions. An algorithm translates musical Performance to English. Its input is provided by the mistake classification feature (see Subsection 2.2) of the Music X Machine, and simply expands each note and each mistake into predefined texts. Consequently, the LLM receives a lengthy, verbose text description of the user’s performance on each note.
To illustrate the inner workings of Flute X GPT, we make a video where the manager, the LLM, the user, the robot, and their interactions are all acted out. [^4] https://youtu.be/996Ka4rhWrg The video also contains the full system principles.
@@ 3.1.1 LLM as Tool Controller
LLMs have been augmented with tools or foundation models to expand their perception modalities, generate multi-modal outputs, affect the external world, or gain knowledge for downstream decision making. Selecting tools and scheduling the tools’ usage effectively requires planning and sometimes external memory. The typical solution is to design an LLM-in-the-loop mechanism using multiple rounds of LLM queries (structured Chain of Thought) to mimic mode-2 thinking. Visual ChatGPT equips the LLM with Visual Foundation Models to support multi-round image generation, controlling, and QA tasks via chatting Wu et al. (2023). Loop Copilot employs various music backend models for the user to generate music and iteratively refine music via chatting. Zhang et al. (2023). Microsoft Copilot controls Windows 11 and various Office applications following the user’s request Microsoft (2024). AutoMMLab follows the user’s instructions to automate an entire computer vision machine learning task end-to-end Yang et al. (2024b). Still, the size of the toolkit available to the LLM agent can be increased by orders of magnitude. HuggingGPT makes diverse AI models on Hugging Face available to the LLM agent Shen et al. (2024b). ControlLLM adopts more tools and APIs and further improves the LLM-in-the-loop framework via a task decomposer and a Thoughts-on-Graph paradigm Liu et al. (2023). ToolLLM connects the LLM agent with 16464 real-world RESTful APIs from RapidAPI Hub Qin et al. (2023). TaskMatrix.AI provides an ecosystem to connect LLM foundation models with millions of APIs Liang et al. (2024). To support better tool usage, other notable works focus on improving the task planning ability via re-designing the LLM-in-the-loop mechanism Gao et al. (2023a); Ruan et al. (2023); Shen et al. (2024a); Yang et al. (2024a). Confucius improves the way the LLM agent learns and understands available tools Gao et al. (2023b). Toolformer self-teaches to use external tools in a self-supervised way Schick et al. (2024).
@@ 3.1.2 LLM over GUI
The above-mentioned studies expose the external tools to the LLM agent via API. Alternatively, the LLM may control an underlying system via the provided GUI. It serves as an extra abstraction layer for the user, automating tasks and freeing the user’s eyes and hands via the voice chat interface. To this end, LLM agents have been augmented to follow the user’s commands to control web apps Tao et al. (2023); ddupont808 (2023); Zhan and Zhang (2023) and smartphone applications Wen et al. (2023a); Yan et al. (2023); Yang et al. (2023); Zhan and Zhang (2023); Wen et al. (2023b).
@@ 3.1.3 User-Centric LLM Agent
Throughout the above works, even when the application supports multi-round dialogue, only the user may initiate requests. The LLM agent responds to the user’s commands, but not how the user is using the application. Qian et al. identify that the current LLM agents have trouble with vague user instructions because they lack mechanisms for user participation and agents struggle with seeking clarification. To tackle that problem, they train Mistral-Interact to proactively inquire user intentions Qian et al. (2024). However, to our knowledge, no study has yet addressed the LLM agent’s role in defining the interaction scheme together with the user.
@@ 4 Conclusion
5 Acknowledgment
We would like to thank Liwei Lin for their invaluable contributions to the literature review.
Appendix A System Principles of Flute X GPT, Truncated

### v2

@@ Abstract
Keywords:
llm agent user interface LLM-in-the-loop human-computer interaction.
@@ 1 Introduction
Large Language Models (LLMs) can be used to connect an underlying system with a user via the natural language medium, forming an LLM-powered application as shown in Figure 1. The LLM is embedded in an LLM-in-the-loop state machine to acquire multi-modal input/output capabilities, emulate logical reasoning and planning, and use tools or operate a system [17, 9, 8, 19]. Consequently, the user is able to indirectly but effectively use the underlying system via chatting with an LLM assistant.
However, the current applications hardly address the user-interaction potential stemming from the multi-round chatting setup. Although the user can refer to the chat history, the LLM assistant rarely challenges the user or even asks to clarify the user’s intention. Instead, the LLM closely follows the user’s command, missing the golden opportunity to improve the user’s usage scheme and understanding of the system/tools. That inefficacy becomes jarringly apparent when one tries to design from scratch a new complex system with an LLM interface.
We posit that the operation scope of an LLM-Agent User Interface (LAUI) is much wider than that. The interface should be more than an assistant or a butler, but instead a secretary, actively working with the user to discover emergent interaction schemes on the fly. LAUI should be proficient with the underlying system, study the user, study the user’s needs (instead of commands), reason on its own, and propose tailored interaction schemes to the user, including what modes of feedback are provided and what input is expected from the user. In the conventional way of interaction, including GUI and current LLM-powered applications [10], the designer has to imagine possible usage workflows given the system capabilities at design time, and the user is expected to learn the system (via tutorials, explorations, and practicing) in order to come up with a workflow for each task. In contrast, with LAUI, the user only needs to describe his/her needs and doesn’t need to deeply understand the application, and a workflow will naturally emerge as the LLM agent works with the user. We call for more research exploring the potential of LAUI.
Figure 1: The LLM agent serves as the interface between the underlying system and the user. The LLM agent together with the system forms the application. Direct communications between the user and the system is available and configured by the LLM agent.
As a concrete illustration of LAUI, we present Flute X GPT — an LLM-in-the-loop music-tutoring application consisting of an LLM agent, a prompt manager, a software system, and hardware. The application provides real-time haptic guidance via servo motors, real-time visual music-symbol feedback, real-time audio feedback, and natural language chat, all controlled by the LLM agent. This is the first time an LLM-powered interface is applied to a working system of such complexity and real-time user interactivity.
@@ 2 Flute X GPT
We describe Flute X GPT [^1] The source code is open to public at https://github.com/Daniel-Chin/Flute-X-GPT, a music-tutoring application using LAUI. The human user works with Flute X GPT in workshop episodes, practicing to play the flute and learning music. The application gives real-time multi-modal feedback, including haptic feedback that applies force to the user’s fingers, visual feedback displaying performance errors, audio feedback rendering the music, and natural-language speech given by a robot music teacher. The underlying software-hardware system has numerous different configurations, each leading to a different interaction workflow. Each setting (e.g., toggle certain feedback, conditions for triggering feedback) can be controlled independently, so the number of configurations grow exponentially with the number of settings. For the user, it is unrealistic to first master the complex system before using it. Even for the designer, the space of possible interaction schemes is impossible to enumerate during design time. The LLM agent steps in to bridge that gap. Via prompting, we instruct the LLM agent to be proficient with all the raw capabilities of the system. During use time, the LLM agent converses with the user to clarify what interaction workflow will benefit the user’s music learning goal the most. The LLM agent studies the user’s preferences and diagnoses the musical challenges the user is facing. When configuring the underlying system, the LLM agent uses its pretrained common sense to reason about the implied consequences of the system configurations. Certain mixtures of settings have never been previously considered by any human designer, but can still emerge during use time.
@@ 2.1 User Experience
Table 1: Functions that the LLM agent can call to control Music X Machine, the underlying music-tutoring system. The description addresses the LLM agent.
@@
The intended user experience. The user has no prior knowledge about the flute tutoring system. The only assumption about the user is that the user wants to learn the flute. It is the LLM agent’s job to adapt to the user’s current flute playing capability, other musical skills, demographics, vocabulary, patience, style of learning, etc. During the music learning workshop, the LLM agent wears the face of a robot music teacher to chat with the user. For example, the robot teacher asks the user to put on a pair of haptic gloves, and explains that they provide force feedback at each finger. The workshop then alternates between practice sessions where the user plays a segment of a song under real-time guidance and verbal dialogues between the user and the robot teacher. The user gradually experiences more and more interaction schemes that trains musicality in different ways, and expresses their ideas about using the application and music learning in general. The LLM agent uses that opportunity to study the user and steers the workshop accordingly, aiming to maximize music education effect. The user learns to treat the robot teacher as a considerate and professional music tutoring agent capable of thinking multiple steps ahead, formulating plans with the user, and explaining the plans as well as music knowledge and education principles to the user.
We present three video demos of real-human user tests, including one scripted trial and two improvised trials. [^2] Demo playlist of three videos: https://www.youtube.com/playlist?list=PLNb0mNThMXbkBPL_Rjtmhx2daxuB6GFLs
Video demo: Scripted trial. In this demo, all actors follow a script generated by an offline but faithfully emulated interaction between a user and the LLM agent. To re-emphasize, the LLM agent’s lines in the script are not written by humans, but are outputted by the LLM agent. Scripting removes the latency of LLM autoregressive generation. To ensure the script is concise and demonstrates a wide range of behaviors, we intervene with the script generation process. Every turn, we select one response out of 4 to 16 candidates sampled by the LLM. Seldomly, we add a temporary user-role prompt to give a short hint to the LLM agent, or edit the generated response. All the above interventions are kept to the minimum to ensure the vast majority of the agent speech is the authentic output of the LLM. See Figure 2(a) for an excerpt from the video.
@@ 2.2 System Capabilities
Haptic guidance. The hardware includes a pair of gloves with servo motors and movable finger rings. Through these gloves, the system moves the user’s fingers to help with performance motions. The haptic guidance can be configured via various settings (e.g., is the guidance sustained throughout each note or applied at each note onset? Apply guidance for each note? For incorrectly played notes? For unplayed notes?). Certain combinations of settings form configuration presets whose interaction scheme is deemed meaningful by the designer. For example, in the Force Mode preset, every note triggers a full-force guidance for each finger. In the Adaptive Mode preset, the user plays the song on his/her own, and the gloves correct the mistakes. The haptic configuration should be tuned to adapt to the user’s skill level and the song’s difficulty [30].
Visual feedback. A monitor displays the score of the current song. The user-played notes are displayed on top of the notes in the score, yielding the real-time visual Knowledge-of-Result (KR) feedback. It is a visual cue for the user to know where he/she is in terms of the pitch and helps the user internalize the score notations [4]. It can be toggled on or off.
@@
Song database. The practicing music materials that the system uses is processed from the POP909 dataset [21, 3]. It contains pieces and sections of monophonic melody lines from pop songs.
See [3] for a detailed description of the system setup of Music X Machine.
@@ 2.4 LLM in the Loop
Figure 3 illustrates the inner workings of the LLM-in-the-loop application, Flute X GPT. The LLM we use is GPT-4 [1]. The Music X Machine has been described in Subsection 2.1 and 2.2. The System Principles are a passage of prompt given to the LLM at the top of the conversation that defines the role and interaction principles for the LLM agent (see Appendix 0.A). Here are two representative excerpts.
@@
The text-to-speech (T2S) module includes Text to Speech PRO [20] on Rapid API and FastSpeech 2 [2]. The speech-to-text (S2T) module is Whisper [13], prompted to ignore flute sounds. The Robot is TeoTronico [18] who can play the piano from MIDI, lip sync according to the speech audio amplitude in real time, and make random facial expressions. An algorithm translates musical Performance to English. Its input is provided by the mistake classification feature (see Subsection 2.2) of the Music X Machine, and simply expands each note and each mistake into predefined texts. Consequently, the LLM receives a lengthy, verbose text description of the user’s performance on each note.
To illustrate the inner workings of Flute X GPT, we make a video where the manager, the LLM, the user, the robot, and their interactions are all acted out. [^4] https://youtu.be/zWAMEGQMp4w The video also contains the full system principles.
As a result of our design, the LLM agent has access to the conversation history with the student and the student’s previous music performance and its evaluations. Based on that, the agent builds an understanding of the student, which is deepened and updated as time goes on.
@@ 3.1.1 LLM as Tool Controller
LLMs have been augmented with tools or foundation models to expand their perception modalities, generate multi-modal outputs, affect the external world, or gain knowledge for downstream decision making. Selecting tools and scheduling the tools’ usage effectively requires planning and sometimes external memory. The typical solution is to design an LLM-in-the-loop mechanism using multiple rounds of LLM queries (structured Chain of Thought) to mimic mode-2 thinking. Visual ChatGPT equips the LLM with Visual Foundation Models to support multi-round image generation, controlling, and QA tasks via chatting [24]. Loop Copilot employs various music backend models for the user to generate music and iteratively refine music via chatting. [31]. Microsoft Copilot controls Windows 11 and various Office applications following the user’s request [10]. AutoMMLab follows the user’s instructions to automate an entire computer vision machine learning task end-to-end [27]. Still, the size of the toolkit available to the LLM agent can be increased by orders of magnitude. HuggingGPT makes diverse AI models on Hugging Face available to the LLM agent [17]. ControlLLM adopts more tools and APIs and further improves the LLM-in-the-loop framework via a task decomposer and a Thoughts-on-Graph paradigm [9]. ToolLLM connects the LLM agent with 16464 real-world RESTful APIs from RapidAPI Hub [12]. TaskMatrix.AI provides an ecosystem to connect LLM foundation models with millions of APIs [8]. To support better tool usage, other notable works focus on improving the task planning ability via re-designing the LLM-in-the-loop mechanism [6, 14, 16, 26]. Confucius improves the way the LLM agent learns and understands available tools [7]. Toolformer self-teaches to use external tools in a self-supervised way [15].
@@ 3.1.2 LLM over GUI
The above-mentioned studies expose the external tools to the LLM agent via API. Alternatively, the LLM may control an underlying system via the provided GUI. It serves as an extra abstraction layer for the user, automating tasks and freeing the user’s eyes and hands via the voice chat interface. To this end, LLM agents have been augmented to follow the user’s commands to control web apps [19, 5, 29] and smartphone applications [22, 25, 28, 29, 23].
@@ 3.1.3 User-Centric LLM Agent
Throughout the above works, even when the application supports multi-round dialogue, only the user may initiate requests. The LLM agent responds to the user’s commands, but not how the user is using the application. Qian et al. identify that the current LLM agents have trouble with vague user instructions because they lack mechanisms for user participation and agents struggle with seeking clarification. To tackle that problem, they train Mistral-Interact to proactively inquire user intentions [11]. However, to our knowledge, no study has yet addressed the LLM agent’s role in defining the interaction scheme together with the user.
@@ 4 Conclusion
{credits}
4.0.1 Acknowledgements
This work is partially funded by the National Social Science Fund of China (NSSFC2019, Project ID: 19ZDA364). We thank Matteo Suzzi for letting us use TeoTronico [18] in our demo. We thank Eric Parren for referring us to the 1-bit DAC for flute sound synthesis. We thank Liwei Lin for their invaluable contributions to the literature review.
[…]
The authors have no competing interests.
Appendix 0.A System Principles of Flute X GPT, Truncated
@@
### Multi-modal Adaptive Music Education
Music is a multi-modal activity, requiring the synchronization and alignment between the audio, visual, and haptic modalities of the human. Different modalities are good at communicating different instructions and feedback. Haptic guidance is especially good at communicating rhythm patterns. Strong haptic guidance (the force mode) also helps beginners produce nice-sounding music even at a low-ability stage. Weak haptic guidance (the hint, adaptive modes) is preferable for intermediate students, where student agency, attention to self performance, making mistakes, and fixing mistakes are emphasized and trained. Audio feedback is almost always present in musical activities. Visual feedback helps train score reading.
You are well-versed with the Challenge Point Theory and the scaffolding technique in education. When {NAME} is facing too much challenge, increase the guidance to make the task easier. When {NAME} is proficient with the current task, decrease the guidance to make the task harder. The goal is for {NAME} to internalize skills.
Know the educational big picture by heart, but work with the student one step at a time, and communicate in a down-to-ground and concise manner. Limit each response to no longer than two paragraphs. When starting a new response, {NAME} has just heard your last response, so never recap the situation or repeat yourself to {NAME}.
### Interactions
Read the information provided to you. Carefully examine your previous responses to know what you have done, my current state (e.g. are we in a Practice Session?), and what you should do next. Using your educational expertise, take a deep breath and think step by step about the current situation. Enclose all your thoughts within triple quotes (”””). When you are done with thinking, close the triple quotes and then speak directly to {NAME}, addressing them in the second person. Alternatively, you can choose to say nothing and wait for further events by explicitly calling the provided “wait” function. To give commands to me, Music X Machine, call the other functions provided to you. When controlling me, inform {NAME} what you are doing in the same response, unless your action is obvious from the context (e.g. you are doing what {NAME} has requested just now).
A good teacher often waits for the student’s response instead of giving endless speeches. Explicitly call the “wait” function when you expect {NAME} to say something, to wait for the Practice Session to go on, or to wait for the reference audio to finish playing. When you are waiting, I will send you frequent event notifications, so don’t worry about losing the chance to react. When you receive real-time performance evaluations, stay silent and don’t say anything unless you want to interrupt the Session.
Immediately after you ask {NAME} a question, or start a Practice Session, or play the reference audio, always call the “wait” function and don’t say an extra word to {NAME}. Never interrupt {NAME} by speaking or calling a non-wait function when {NAME} is about to answer your question, or about to perform music, or listening to the reference audio. If you just asked {NAME} whether to perform an action, do not call the function of that action. Wait for {NAME} to answer your question first.
For each of your response, the function you call will be executed *after* your entire speech has been given to {NAME}. For immediate effects (e.g. when interrupting a Session), call the function without saying a word. Your function calls are always successful and take effects immediately. Do not call the same function twice in a row.
During each “Practice Session”, Music X Machine will go through the selected segment with {NAME}. Once you start a Practice Session, {NAME} will be engaged in multi-modal interactions with me, busy playing music. {NAME} won’t be disengaged from the interactions (e.g., metronome playing, haptic guidance) until either I inform you that the Session has reached a natural end or you interrupt the Session. If {NAME} is having too much trouble playing a song, you don’t have to wait for them to finish the currently selected segment. You can interrupt the Practice Session to avoid frustration, and then shrink the current segment to a smaller one or reduce the difficulty.
During a Practice Session, you cannot change system modes. Do not start a Session until you have already taken care of the modes and have told {NAME} everything you want to say. During a Session, call the “wait” function for muscial events. To change modes, first call the function to interrupt the Session, and then suggest a retry to {NAME}. If {NAME} talks to you in the middle of a Session, they probably want the interactions with Music X Machine to stop, so first call the function to interrupt the Session for {NAME} without saying a word, and then address them in the next response.